A certain area of the eastern United States is, on average, hit by 6 hurricanes a year. Find the probability that in a given year that area will be hit by



(a) fewer than 4 hurricanes;



(b) anywhere from 6 to 8 hurricanes.

1
Expert's answer
2021-11-28T19:00:23-0500

(a) fewer than 4 hurricanes;

"P(X)=\frac{\lambda^{X}e^{-\lambda}}{X!}"


"P(X<4)=P(X=0)+P(X=1)+P(X=2)+P(X=3)"


"=\frac{6^{0}e^{-6}}{0!}+\frac{6^{1}e^{-6}}{1!}+\frac{6^{2}e^{-6}}{2!}+\frac{6^{3}e^{-6}}{3!}"


"=0.0025+0.0149+0.0446+0.0892"

"=0.1512"


(b) anywhere from 6 to 8 hurricanes;

"P(X)=\frac{\lambda^{X}e^{-\lambda}}{X!}"


"P(6\eqslantless X\eqslantless8)=P(X=6)+P(X=7)+P(X=8)"


"=\frac{6^{6}e^{-6}}{6!}+\frac{6^{7}e^{-6}}{7!}+\frac{6^{8}e^{-6}}{8!}"


"=0.1606+0.1376+0.1033"

"=0.4015"